OpenStudy (anonymous):

A triangle has side lengths of 10, 15, and 7. Is the triangle acute, obtuse or right? Explain.

OpenStudy (anonymous):

c^2 = a^2 + b^2 15^2 = 7^2 + 10^2 225 = 49 + 100 225 > 149 Since c^2 > a^2 + b^2, it is obtuse. Here are the rules: c^2 = a^2 + b^2: Right Triangle c^2 > a^2 + b^2: Obtuse Triangle c^2 < a^2 + b^2 Acute Triangle